OpenStudy (jessiegonzales):

medal What is an equation of the line, in point-slope form, that passes through the given points and has the given slope? pont:(8,-8);slope:3 y-8=3(x+8) y+8=3(x+8) y-8=3(x-8) y+8=3(x-8)

OpenStudy (anonymous):

this one easier than the last, since they gave you the slope to being with

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] with \[x_1=8,y_1=-8\] remember the sign comes with the number

OpenStudy (anonymous):

you got it yet?

OpenStudy (jessiegonzales):

ok so a is the answer

OpenStudy (anonymous):

no are you getting the x's confused with the y's?

OpenStudy (jessiegonzales):

i think so

OpenStudy (anonymous):

i can see it is a bit confusing because they use the point \((8,-8)\) plug in \(x\) for \(x_1\) and \(-8\) for \(y_1\) in \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

oops i meant "plug in \(8\) for \(x\)"

OpenStudy (anonymous):

remember that x comes before y in the alphabet, just like it does in \((x,y)\) course you have to sing that whole song until the end

OpenStudy (jessiegonzales):

ok im los t

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

the "point slope" formula for a line, with slope \(m\) and point \((x_1,y_1)\) is \[\huge y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

\(x\) and \(y\) are variables, they stay in the answer \(x_1\) and \(y_1\) are numbers

OpenStudy (anonymous):

you have in this case \[x_1=8,y_1=-8,m=3\] so we plug them directly in to the formula, remembering that the minus sign that is in front of the number stays there

OpenStudy (anonymous):

i will plug them directly in :\[y-(-8)=3(x-8)\]

OpenStudy (anonymous):

then you see that \(-(-8)=+8\) so in the end you get \[y+8=3(x-8)\]

OpenStudy (jessiegonzales):

now i see thanks